What length of copper wire 1 mm in diameter is needed to make a 30 ohm resistor at 18 degrees C.
Answer #1

the resistivity of Copper at 20 °C is 1.68 * 10-8 , it's temperature coefficient is 0.0039 ℃-1

∆T = 18 -20 = -2℃

= 1.68×10-8×(1+0.0039×(-2))

= 1.667×10-8

Cross section area is

A = πd2/4

= π ×(1×10-3 m)2/4

= 0.785 ×10-6 m2

= 30×0.785×10-6/(1.667×10-8)

=  1413 m